find the domain and range of: A)y=3x-9 B)y=-|x|+2 C)y=2x^2+6x-7
Answer by stanbon(75887)

find the domain and range of:
A)y=3x-9
Domain and Range are "All Real Numbers".
---------------------------------------
B)y=-|x|+2
Domain: All Real Numbers
Range: All Real Numbers <= 2
===============
C)y=2x^2+6x-7
Find the vertex:
x = -b/(2a) = -6/(-4) = 3/4
Then f(3/4) = -30.38
----
Domain: All Real Numbers
Range:: All Real Numbers >= -4.061
=======================
Cheers,
Stan H.
==========